Consider a bond that has a 30-year maturity, an 8% coupon rate, and sells at an...

  1. Consider a bond that has a 30-year maturity, an 8% coupon rate, and sells at an initial yield to maturity of 8%. Because the coupon rate equals the yield to maturity, the bond sells at par value: P = $1,000.00. Calculate the duration and the modified duration. If we assume the convexity of the bond is 212.4 and the bond’s yield increases from 8% to 10%, how much should the bond price decline?
Answer #1

Period Cash Flow Discounting factor PV Cash Flow Duration Calc
0 ($1,000.00) =(1+YTM/number of coupon payments in the year)^period =cashflow/discounting factor =PV cashflow*period
1             80.00                                                             1.08                    74.07                  74.07
2             80.00                                                             1.17                    68.59                137.17
3             80.00                                                             1.26                    63.51                190.52
4             80.00                                                             1.36                    58.80                235.21
5             80.00                                                             1.47                    54.45                272.23
6             80.00                                                             1.59                    50.41                302.48
7             80.00                                                             1.71                    46.68                326.75
8             80.00                                                             1.85                    43.22                345.77
9             80.00                                                             2.00                    40.02                360.18
10             80.00                                                             2.16                    37.06                370.55
11             80.00                                                             2.33                    34.31                377.42
12             80.00                                                             2.52                    31.77                381.23
13             80.00                                                             2.72                    29.42                382.41
14             80.00                                                             2.94                    27.24                381.32
15             80.00                                                             3.17                    25.22                378.29
16             80.00                                                             3.43                    23.35                373.62
17             80.00                                                             3.70                    21.62                367.57
18             80.00                                                             4.00                    20.02                360.36
19             80.00                                                             4.32                    18.54                352.20
20             80.00                                                             4.66                    17.16                343.28
21             80.00                                                             5.03                    15.89                333.74
22             80.00                                                             5.44                    14.72                323.74
23             80.00                                                             5.87                    13.63                313.38
24             80.00                                                             6.34                    12.62                302.78
25             80.00                                                             6.85                    11.68                292.04
26             80.00                                                             7.40                    10.82                281.22
27             80.00                                                             7.99                    10.01                270.40
28             80.00                                                             8.63                      9.27                259.65
29             80.00                                                             9.32                      8.59                249.00
30       1,080.00                                                          10.06                  107.33              3,219.83
      Total            12,158.41
Macaulay duration =(∑ Duration calc)/(bond price*number of coupon per year)
=12158.41/(1000*1)
=12.158406
Modified duration = Macaulay duration/(1+YTM)
=12.16/(1+0.08)
=11.257783
New bond price @ YTM =10 using duration
Mod.duration prediction = -Mod. Duration*Yield_Change*Bond_Price
=-11.2577*0.02*1000
=-225.154
New bond price = bond price+Modified duration prediction
=1000+-225.154
=774.85
New bond price @ YTM =10 using duration and convexity
Convexity adjustment = 0.5*convexity*Yield_Change^2*Bond_Price
=0.5*212.4*0.02^2*1000
=42.48
New bond price = bond price+Mod.duration pred.+convex. Adj.
=1000+-225.15+21.24
=817.33

Price declined by 1000-817.33=182.67